OpenStudy (anonymous):

for systems of linear equations in two variables, 2x+y=4 x-3y=9 do i just graph and find where the lines meet?

OpenStudy (anonymous):

2X+y=4 X-3y=9 first get rid of x or y on both equation

OpenStudy (anonymous):

well what i did was turn it into slope intercept so now they are y=4-2x and y=-3-x

OpenStudy (anonymous):

no look it as if it was an addition or subtraction problem, im drawing a pic its hard to explain. give me a sec

OpenStudy (anonymous):

okay but the instructions say to solve each stem . if the system has an infinite solution set, specify it. if the system has no soulution say so.

OpenStudy (anonymous):



OpenStudy (anonymous):

multiply everythin the first equation by 3

OpenStudy (anonymous):

you should get 6X+3y=12

OpenStudy (anonymous):

how do you know to do that instead of graphing?

OpenStudy (anonymous):

this way you are solving for X and Y, so you know exactly where the two lines meet on a graph

OpenStudy (anonymous):

okay let me solve it real fast and i will tell you with what i end up with

OpenStudy (anonymous):

so far i have x=3 then i plug it into one of the original problems right?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

okay my final answer turned out to be (3,-2) thank you so much!!

OpenStudy (anonymous):

hold on give me a sec sorry,

OpenStudy (anonymous):

actually y should be 14

OpenStudy (anonymous):

X is 3/5

OpenStudy (anonymous):

remember X is your slope so its ok to have a fraction Y=mx+b

OpenStudy (anonymous):

this is how i solved it 3(2x+y=4) x- 3y=9 6x+3y=12 7x+21 x+3

OpenStudy (anonymous):

multiply the top by a -3 that way y turns to -3y and you can cancel out the other +3y in the second equation so you are left with -6x=-12 x = 9

OpenStudy (anonymous):

well wouldnt it work if i multiplied the top by 3 in order to cancel out the y coefficiant?